OpenStudy (anonymous):

help me i need to pass my test PLease The first term of a geometric series is 3, and the sum of the first term and the second term is 15. What is the sum of the first six terms?

OpenStudy (kropot72):

The first term is 3. The second term is 3 multiplied by the common ratio r. So we can write an equation from the given value of the sum of the first and second terms: 15 = 3 + 3r This equation can be solved to find the value of r. Can you do that?

OpenStudy (anonymous):

r=4

OpenStudy (anonymous):

What about terms 3-6?

OpenStudy (kropot72):

Good work! Now that you have the value of r you can use the following formula to find the sum of the first six terms: \[S _{n}=\frac{a(r ^{n}-1)}{r-1}\]

OpenStudy (anonymous):

28.2

OpenStudy (anonymous):

a.1023 b.3906 C4096 d.11718 these are the options

OpenStudy (anonymous):

2457

OpenStudy (anonymous):

but its the choses i am given

OpenStudy (kropot72):

Sorry, my bad. Substituting in the formula for the sum of n terms gives the following: \[S _{6}=\frac{3(4^{6}-1)}{4-1}=4^{6}-1=4095\]

OpenStudy (anonymous):

thabks

OpenStudy (kropot72):

You're welcome :)